OpenStudy (anonymous):

Pablo is making ramps of different lengths. The ramps are made with rectangular and square wooden boards. The relationship between the length of the ramp in feet, f(n), and the number of square boards used, n, is represented by the function below. f(n) = 20n + 3 , where 2 ≤ n ≤ 6 Which statement is correct about the range of the length of the ramp for the given domain? It varies from 2 to 6. It varies from 43 to 123. It varies from 40 to 120. It varies from 2 to 123. My answer is It varies from 2 to 6. Am I right? :)

jimthompson5910 (jim_thompson5910):

f(n) = 20n + 3 f(2) = 20(2) + 3 f(2) = 40 + 3 f(2) = 43 The smallest output possible is 43

jimthompson5910 (jim_thompson5910):

To find the largest output possible, plug in the largest input possible

OpenStudy (anonymous):

which would be 123?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

smallest output possible: 43 largest output possible: 123 So the range is 43 to 123

OpenStudy (anonymous):

It varies from 40 to 120.